The sum of all 3 digits of a number is identical with the first two digits, and the sum of the sum is identical with the third. What is the number?
Answer by ankor@dixie-net.com(22740)

The three digits, x, y, z
:
The sum of all 3 digits of a number is identical with the first two digits,
x + y + z = 10x + y
z = 10x - x + y - y
z = 9x, therefore
z = 9 and x = 1, (in order for them to to be single digits)
:
and the sum of the sum is identical with the third.
x + y = z
Replace x and z, find y
1 + y = 9
y = 9 - 1
y = 8
:
What is the number? 189 is the number